A gas that evolved during the fermentation of glucose has a volume of 0.78 L when measured at 20.1 °C and 1.0 atm. What was the volume of this gas at the fermentation temperature of 36.5 °C? Assume ideal-gas behavior.
1
Expert's answer
2021-10-18T03:10:42-0400

P1V1/T1= P2V2/T2

1×0.78/293.1= 1×V2/309.5

0.0027= V2/309.5

V2= 0.0027×309.5

V2= 0.84L